OpenStudy (anonymous):

The midpoint of the coordinates (7,9) and (12,11) is ______ a.(9.5,10) b.(13,12) c.(19,20) d.none of the above

OpenStudy (zepp):

Do you know what is the midpoint formula?

OpenStudy (anonymous):

yes y1-y2 x1-x2

OpenStudy (zepp):

That's the formula to find to slope, y2 - y1 / x2 - y1 The midpoint formula is very simple

OpenStudy (zepp):

It's almost the same thing than the slope-point formula. \[(\large \frac{x_{1} + x_{2}}{2}, \frac{y_{1} + y_{2}}{2})\] Basically, you find the 'average' if you want. :)

OpenStudy (anonymous):

i got 19/2,10

OpenStudy (zepp):

Yup, that's right :)

OpenStudy (anonymous):

i don't know if its right.

OpenStudy (zepp):

Great work ;D

OpenStudy (anonymous):

so its d?

OpenStudy (anonymous):

Thanks So Much Friend=)

OpenStudy (zepp):

\[\frac{19}{2}=9.5\]

OpenStudy (zepp):

\[(9.5,10)\] It should be A) You are welcome :)

OpenStudy (amistre64):

im glad im wrong every now and then :) i came to see someone spitting out answers and found a good response ...